- February 9th: Routing issues that were caused by a faulty FPC. We replaced the FPC.
- December 15th: Fiber Loss high alarm observed in UPA-RM1. Alarm cleared after provider disabled and then enabled the fiber detect function.
- February 28th: SUNET upgraded the software version at one of the optical components in Lund.
- February 19th: Duplicate, see SUNET-191.
- February 8th: The SCU in shelf 2 was rebooted by remove and insert in the shelf, the alarms (Mod led was flashing yellow) cleared after his and the node went back to normal. No outage observed.
- February 7th: SUNET performed a maintenance to optimize the optical channel on the link between Kalmar and Kalskrona.
|Hardware||0 pcs||1 pcs||1 pcs|
|Link||0 pcs||1 pcs||1 pcs|
|Other||0 pcs||0 pcs||0 pcs|
|Routing||1 pcs||1 pcs 00:28||2 pcs 00:28|
|Software||0 pcs||1 pcs 00:28||1 pcs 00:28|